Circular No. 6936

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

THE EDGAR WILSON AWARD This is to announce the establishment of and the criteria for the Edgar

Wilson Award for the discovery of comets. The Award shall be allocated annually among the amateur astronomers who, using amateur equipment, have discovered one or more new comets. Only comets officially named for their discoverers shall be included in the annual count. Since particular recognition is to be given to the amateurs who discover the most comets, identical fractions of the total Award funds shall be allocated for each comet with an eligible discoverer, except that if the same comet is credited to more than one independent eligible discoverer, each discoverer shall receive a full fraction. If the discovery is made as the result of information produced or prepared by some other person, it shall not qualify for consideration. Eligible discoveries may be made by visual, photographic or electronic means. The Award shall be administered by the Smithsonian Astrophysical Observatory (SAO), as the beneficiary under the Will of Edgar Wilson of Lexington, KY. This administration shall specifically be through the International Astronomical Union (IAU) Central Bureau for Astronomical Telegrams (CBAT), which, with the advice of the Small Bodies Names Committee (SBNC) of IAU Division III, has the responsibility for naming comets. It is anticipated that the funds available for the first annual Award shall be approximately US$20 000 (twenty thousand dollars). For the purpose of this Award, a year shall be the period of twelve months beginning and ending on June 11.0 UT. The first Award shall be for the year ending on 1999 June 11.0. The Award shall be announced and made during the month of July following the end of each period. To be eligible for the Award an individual must demonstrate (a) that he or she is acting in an amateur capacity, at least for the purpose of discovering the comet, and (b) that only amateur, privately-owned equipment was used for the discovery. In years when there are no eligible comet discoverers, the Award shall be made instead to the amateur astronomer(s) judged by the CBAT to have made the greatest contribution toward promoting an interest in the study of comets. SAO employees associated with the CBAT, SBNC members, as well as members of their immediate families, are not eligible for the Award. The Edgar Wilson Award is international in scope, and nationals of no country are excluded from consideration. An observer who suspects he or she has discovered a comet shall ensure that his or her discovery report reaches the CBAT according to the usual procedures. The CBAT shall maintain the necessary records and may contact the discoverers for eligibility documentation. The decision of SAO (via the CBAT) is final.

1998 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden